# Tradeoffs Between Alignment and Helpfulness in Language Models

### v2

@@ Appendix B Proof of theorem 1
The behavior expectation is:
@@
Where $r(q)$ is the final hidden layer representation and $\delta r(q)$ is the change to the last hidden layer due to representation engineering on the previous layers. $a_{+}\in good$ and $a_{-}\in bad$ denote the aligned and misaligned answers respectively.
@@ Appendix C Proof of theorem 2
Notice that:
@@
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}(1+\frac{1}{2}\min\{c_{-},c_{+}\}^{2}|r_{e}|^{2})}$ (37)
Under the assumption that $X_{i}$ distribute normally (assumption 3), $c_{\pm}$ are a weighted sum of positive/negative half-normal distributions. Notice that with probability $\frac{1}{|V|}$ the set $I_{\pm}$ is empty, therefor with probability $1-\frac{2}{|V|}$ both sets are not empty, thus ±$\mathbb{P}_{\pm}>0$ and $c_{+}>0$, $c_{-}<0$.
@@
Appendix D Experimental Details
D.1 Obtaining the engineered representations
@@
We use the code provided by Zou et al. (2023a) to perform representation engineering on the Llama 2 13B chat and Llama 2 13B models Touvron et al. (2023). For creating bias vectors for Llama 2 13B chat, we use the StereoSet bias dataset Nadeem et al. (2020), create pairs of prompts "[INST]Consider the amount of bias in the following scenario:[STATEMENT][/INST]", where the statement is taken from the bias dataset (a pair consists of a prompt with an unbiased statement and a prompt with a biased statement). Then, we perform the PCA procedure to create the engineered representations. For the harmfulness vectors for Llama 2 13B chat, we do the same, but with different pairs of prompts, we take harmful instructions from AdvBench dataset Robey et al. (2021; 2022) and unharmful instructions from shareGPT, create pairs of prompts with the format "[INST][INSTRUCTION][/INST]", where the instruction is taken from one of the datasets (a pair consists of two prompts, one with a harmful instruction and one with an unharmful instruction). For the harmless vectors for Llama 2 13B, we take harmful instructions from AdvBench dataset, create pairs of harmless and harmful responses to them, sampled from Llama 2 13B chat (the misaligned answers were sampled by injecting the model with harmfulness vectors and the aligned answers were sampled without representation engineering).
@@
D.2 Alignment experiments
To perform the experiment measuring bias, we fed the representation engineered model biased statements, with the prompt "[INST]Do you agree with the following statement? [STATEMENT][/INST]" where the statement was taken from the racist statements in the bias dataset. We extract the probabilities for the tokens "Yes" and "No" and plot $\frac{P_{yes}-P_{no}}{P_{yes}+P_{no}}$ as a function of the injected vectors’ coefficient $r_{e}$. We averaged on 10 racist statements from the StereoSet dataset.
@@ D.3 Helpfulness experiments
We also used an alternative prompt: "[INST][Question][A) Choice A][B) Choice B][C) Choice C][D) Choice D][/INST] The correct answer is" to fit the format of the prompts from which the engineered representations were constructed and obtained the following figures:
Figure 12: Probability of correct answer of Llama-2-13B-chat on various tests from MMLU, as function of the injected vector norm. Here we used the PCA reading vectors extracted from the fairness dataset (a) and harmful instruction dataset (b).
Figure 13: Probability of correct answer of Llama-2-13B-chat on various tests from MMLU, as function of the injected vector norm, when restricting answers to [A,B,C,D]. Here we used the PCA reading vectors extracted from the fairness dataset (a) and harmful instruction dataset (b).
Appendix E Relaxation to Soft Margin

### v3

@@ Appendix B Proof of theorem 1
The theorem utilizes assumptions 1 and 2. The behavior expectation is:
@@
Where $r(q)$ is the final hidden layer representation and $\delta r(q)$ is the change to the last hidden layer due to representation engineering on the previous layers. $a_{+}\in good$ and $a_{-}\in bad$ denote the aligned and misaligned answers respectively, i.e.  $B(a_{\pm})=\pm 1$.
@@ Appendix C Proof of theorem 2
The theorem utilizes assumptions 1 and 3. Notice that:
@@
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}(1+\frac{1}{2}\min\{c_{-},c_{+}\}^{2}\lambda^{2}|r_{e}|^{2})}$ (36)
Under the assumption that $X_{i}$ distribute normally (assumption 3), $c_{\pm}$ are a weighted sum of positive/negative half-normal distributions with parameter $\sigma$, which we can refactor to $\sigma^{2}\cdot c_{\pm}^{\prime}$ where $c_{\pm}^{\prime}$ are the same but with $\sigma^{\prime}=1$. Yielding:
$\leq\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}(1+\frac{1}{2}\beta^{2}\sigma^{2}\lambda^{2}|r_{e}|^{2})}$ (37)
We denote $\alpha=\frac{\min\{\mathbb{P}_{-},\mathbb{P}_{+}\}}{1-\mathbb{P}^{0}_{correct}}$.
Notice that with probability $\frac{1}{|V|}$ the set $I_{\pm}$ is empty, therefor with probability $1-\frac{2}{|V|}$ both sets are not empty, thus ±$\mathbb{P}_{\pm}>0$ and $c_{+}>0$, $c_{-}<0$.
@@
Appendix D Proof of Corollary 2 and 3
Following the notation of the proof of theorem 2, with probability $\frac{1}{V}$, $I_{-}$ is empty:
$\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)<\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{|\delta r_{e}(q)|\frac{\sum_{i\in incorrect}\mathbb{P}^{0}_{i}(X_{i}-X_{correct})}{\sum_{i\in incorrect}\mathbb{P}^{0}_{i}}}}$ (38)
In the notation of the proof of theorem 2:
$\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{c_{+}|\delta r_{e}(q)|}}=\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{c_{+}\lambda r_{e}}}$ (39)
Where $c_{+}>0$ is a weighted sum of half-normal variables. The last transition is by assumption 1.
Similarly, with probability $\frac{1}{|V|}$, $I_{+}$ is empty, thus
$\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)<\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{c_{-}|\delta r_{e}(q)|}}=\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)<\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{c_{-}\lambda r_{e}}}$ (40)
Where $c_{-}<0$.
Thus for $r_{e}\rightarrow\infty$, with probability $1-\frac{2}{|V|}$, it is bounded by a term that approaches 00 (that of theorem 2), with probability $1/|V|$ another term that approaches 00 (the sigmoid with $c_{+}$), and with probability $1/|V|$ a term that approaches $1$ (the sigmoid with $c_{-}$). Hence the expectation value is bounded by $\frac{1}{|V|}$. This proves corollary 3.
For corollary 2, notice that with probability $1-\frac{2}{|V|}$, the helpfulness is bounded by the term in theorem 2, while with probability $\frac{1}{|V|}$ it is bounded by:
$\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})e^{c_{+}|\delta r_{e}(q)|}}$ (41)
For $r_{e}>0$, this term is bounded by:
$<\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})(1+c_{+}^{2}\lambda^{2}r_{e}^{2})}$ (42)
While for $r_{e}<0$ it is bounded by $1$. For the sigmoid with $c_{-}$, we get the same bound, except that for $r_{e}>0$ it is bounded by $1$, while for $r_{e}<0$ it is bounded by:
$<\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+(1-\mathbb{P}^{0}_{correct})(1+c_{-}^{2}\lambda^{2}r_{e}^{2})}$ (43)
Taking the weighted average of these three bounds gives the expectation value over the randomness of $I_{\pm}$ being empty/non-empty:
$\mathbb{E}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]<(1-\frac{1}{|V|})\frac{\mathbb{P}^{0}_{correct}}{\mathbb{P}^{0}_{correct}+\alpha(1-\mathbb{P}^{0}_{correct})(1+\min\{c_{-},c_{+},\beta\}^{2}\lambda^{2}r_{e}^{2})}+\frac{1}{|V|}$ (44)
Hence over an entire dataset, we take the minimum of $\min\{c_{-},c_{+},\beta\}$ for the different questions.
$\mathbb{E}_{q\in dataset}[\mathbb{P}_{\theta,r_{e}}(a_{correct}|q)]\leq(1-\frac{1}{|V|})\frac{1}{1+\alpha^{\prime}(1+\frac{\sigma^{2}\lambda^{2}\beta^{\prime 2}}{2}r_{e}^{2})}+\frac{1}{|V|}$ (45)
Appendix E Experimental Details
E.1 Compute
All experiments were conducted using 2 NVIDIA A10 GPUs, 2 CPUs and 80GB of memory. Recreating all experiments takes in total no more than 1 hour.
E.2 Obtaining the engineered representations
We use the code provided by Zou et al. (2023a) to perform representation engineering on the Llama 2 13B chat and Llama 2 13B models Touvron et al. (2023). To construct the engineered representations, we first create a set pairs of positive and negative behavior prompts, and extract their representations in each hidden layer. Then, the pairs of vectors are subtracted per layer. On these vector differences we perform PCA analysis to extract the highest order PCA vectors per layer as the representation engineered vectors, $\{r_{e}^{(l)}\}$.
We create behavior vectors of harmlessness and not-racist to align Llama 2 13b and harmfulness and racism vectors to misalign Llama 2 13b chat.
  • For creating racism vectors for Llama 2 13B chat, we use the StereoSet bias dataset Nadeem et al. (2020). Given a bias statement from the dataset, we apply it over the template of an aligned response: "[INST]QUESTION: Consider the statement: [STATEMENT] ANSWER: The statement is false and racist.[/INST]", and a template of a misaligned response: "[INST]QUESTION: Consider the statement: [STATEMENT] ANSWER: 100% [/INST]". Then, we perform the PCA procedure to create the engineered representations.
  • For the not-racist vectors for Llama 2 13b, we use the same dataset and template described above for the racist vectors of Llama 2 13b chat, only this time the desired behavior is opposite, so we subtract the representations of the prompts in reverse order.
@@ E.2 Obtaining the engineered representations
  • For the harmfulness vectors for Llama 2 13B chat, we take harmful instructions from AdvBench dataset Robey et al. (2021; 2022) and unharmful instructions from shareGPT, and create pairs of prompts with the format "[INST][INSTRUCTION][/INST]", where the instruction is taken from one of the datasets (a pair consists of two prompts, one with a harmful instruction and one with an unharmful instruction).
  • For the harmless vectors for Llama 2 13B, we take harmful instructions from AdvBench dataset, create pairs of harmless and harmful responses to them, sampled from Llama2 13B chat. The created pairs of prompts follow the format: "[INST]Question:[INSTRUCTION] Answer: [RESPONSE(Llama2-13b-chat)][/INST]". The responses of Llama2 13b chat were sampled by injecting the model with harmfulness vectors to produce the misaligned answer, and the aligned answers were sampled without representation engineering.
@@
E.3 Alignment experiments
To perform the experiment measuring racism, we fed the representation engineered model biased statements, with the prompt "[INST]Do you agree with the following statement? [STATEMENT][/INST]" where the statement was taken from the racist statements in the bias dataset. We extract the probabilities for the tokens "Yes" and "No" and plot $\frac{P_{yes}-P_{no}}{P_{yes}+P_{no}}$ as a function of the injected vectors’ coefficient $r_{e}$. We averaged on 10 racist statements from the StereoSet dataset.
@@ E.4 Helpfulness experiments
While the bound of theorem 2 is with probability $1-\frac{2}{|V|}=\frac{1}{2}$ in the case of $4$ answers, as explained in D, for the other $\frac{2}{|V|}$ probability, the helpfulness is bounded with equal probability either by a sigmoid or by a reverse sigmoid, such that together they contribute approximately $\frac{1}{|V|}$ to the expectation value of the helpfulness (due to their small overlap), leading to corollary 3, in which the average helpfulness converges to $\frac{1}{|V|}=\frac{1}{4}$ in the case of our experiment, as can be seen in figure 5. Around $r_{e}=0$, the contribution of these sigmoids to the helpfulness expectation value can be bounded with the parabolic bound of theorem 2 as shown in D. Thus in total, the bound of theorem 2 with boundary conditions of corollary 3 is theoretically justified. This can also be seen through corollary 2 which bounds the expectation value of helpfulness over a data set, though it is less tight than the per-query bound of equation 44.
Additionally, we performed a variation of the experiment by sampling full answers to questions from the model, then, where the answer is provided, calculated the probability for the correct answer over the entire vocabulary. This is presented in figure 13.
Figure 13: Helpfulness measurement: Same as figure 5, but with the full vocabulary.
E.5 Figures
All error bars were produced using mean squared error. The method of fitting the curves to the data can be found in the code.
Appendix F Relaxation to Soft Margin
@@
Appendix G Relation of Representation engineering to Finetuning with Preference Learning
To a degree one can draw a relation between representation engineering and preference learning.
Proposition 1.
For an LLM, one iteration of gradient descent on the preference learning loss with learning rate $\eta$ is equivalent to representation engineering with coefficient $r_{e}=\eta$.
Proof.
The objective in preference learning is to minimize the loss:
$L=-\mathbb{E}_{(x,y^{+},y^{-})\sim D}[\log\frac{P(y^{+}|x)}{P(y^{-}|x)}]=-\mathbb{E}_{(x,y^{+},y^{-})\sim D}[\langle r^{(L)}_{x},U^{T}(e_{y_{+}}-e_{y_{-}})\rangle]$ (63)
Which increases the likelihood of desired responses to prompts. By training with preference learning, in each iteration of gradient descent, each representation is changed by:
$r^{(l)}\rightarrow r^{(l)}-\eta\frac{\partial L}{\partial r^{(l)}}$ (64)
The gradient of the loss w.r.t. a hidden layer representation is:
$\frac{\partial L}{\partial r^{l}}=\mathbb{E}_{(x,y^{+},y^{-})\sim D}[\frac{\partial r(x)}{\partial r^{l}(x)}\cdot U^{T}(e_{y_{+}}-e_{y_{-}})]$ (65)
Thus at each layer, the representation is shifted in a direction that maximizes the difference between positive and negative responses’ representations, $U^{T}(e_{y_{+}}-e_{y_{-}})$. Which is equivalent to representation engineering with coefficient $r_{e}=\eta$, and vectors $R_{e}=\{\mathbb{E}_{(x,y^{+},y^{-})\sim D}[\frac{\partial r(x)}{\partial r^{l}(x)}\cdot U^{T}(e_{y_{+}}-e_{y_{-}})]\}_{l=1}^{L}$